[C] Solve the following differential equation:

x^ 2 y^ " - 2xy'-4y=x^ 2 +2 log x


1
Expert's answer
2021-12-20T19:03:00-0500

Let us solve the differential equation "x^ 2 y''- 2xy'-4y=x^ 2 +2 \log x."

Let us use the transformation "x=e^t." Then "y'_x=y_t'e^{-t},\ y_{x^2}''= (y_{t^2}''-y_t')e^{-2t}."

We get the equation "e^{2t}(y_{t^2}''-y_t')e^{-2t}- 2e^{t}y_t'e^{-t}-4y=e^{2t} +2t," which is equivalent to

"y_{t^2}''- 3y_t'-4y=e^{2t} +2t."

The characteristic equation "k^2-3k-4=0" of the differential equation "y_{t^2}''- 3y_t'-4y=0" is equivalent to"(k-4)(k+1)=0," and hence has the roots "k_1=4,\ k_2=-1."

Therefore, the general solution of the last differential equation is "y(t)=C_1e^{4t}+C_2e^{-t}+y_p," where "y_p=ae^{2t}+bt+c." It follows that "y_p'=2ae^{2t}+b,\ y_p''=4ae^{2t}," and hence we get

"4ae^{2t}-3(2ae^{2t}+b)-4(ae^{2t}+bt+c)= e^{2t} +2t."

It follows that "-6ae^{2t}-4bt-3b-4c= e^{2t} +2t," and hence "-6a=-1,\ -4b=2,\ -3b-4c=0." Consequently, "a=\frac{1}6,\ b=-\frac{1}2,\ c=-\frac{3}4b=\frac{3}{8}."

We conclude that "y(t)=C_1e^{4t}+C_2e^{-t}+\frac{1}6e^{2t}-\frac{1}2t+\frac{3}8."


Therefore, the general solution of the differential equation "x^ 2 y''- 2xy'-4y=x^ 2 +2 \log x" is "y(x)=C_1x^4+C_2x^{-1}+\frac{1}6x^2-\frac{1}2\log x+\frac{3}8."